DIFFERENTIAL GEOMETRY AND APPLICATIONS Satellite Conference of ICM in Berlin, Aug. 10 – 14, 1998, Brno Masaryk University in Brno (Czech Republic) 1999, 285–292

NATURAL OPERATORS ON THE

BUNDLE OF CARTAN CONNECTIONS

Martin Panák

Abstract. We show that Cartan connections of a given type can be viewed as sections of a gauge natural bundle. This fact enables us to study natural operators on Cartan connections as gauge natural invariants. We deduce a modification of the Utiyama theorem. The relation to the classical results on natural operators on affine connections is discussed in Section 7.

1. Cartan connections

A Cartan connection on a principal B-bundle P over a m-dimensional manifold is a 𝔤-valued one-form ω (𝔤 is the Lie algebra of G, B is a closed subgroup of G) with the following properties:

(1) ω(X^{∗}) = X, for X ∈ 𝔟, where X^{∗} is the fundamental vector field corresponding to X

(2) r_{a}^{∗}ω = Ad(a^{−1})ω for each a ∈ B, where r_{a} is the right principal action of an element a,

(3) ω : T_{u}P → 𝔤 is an isomorphism for each u ∈ P. We point out, that dim G/B = m.zech

Associating a group G via the left multiplication to a principal B-bundle we get a G-principal bundle P′ = P ×_{B} G with the right principal action (u ×_{B} g) · h = (u ×_{B} g · h). Thus the category PB_{m}(B), the category of principal bundles with structure group B, can be seen as a ”subcategory” of PB_{m}(G), the category of principal bundles with a structure group G. Namely each principal B-bundle P can be identified with a subbundle of P ×_{B} G (u ↦ (u×_{B} e)), and a PB_{m}(B)-morphism φ : P → R yields a morphism φ′ : P′ → R′, φ′(u ×_{B} g) = (φ(u) ×_{B} g). Further we can identify (ℝ^{m} × B) ×_{B} G with ℝ^{m} × G ((x, b) ×_{B} g ↦ (x, bg)). Notice that the composition of the embeddings described above, ℝ^{m} × B → (ℝ^{m} × B) ×_{B} G → ℝ^{m} × G, is just the canonical embedding ℝ^{m} × B → ℝ^{m} × G.

1991 Mathematics Subject Classification. 58A20, 53A55.

Key words and phrases. Cartan connection, natural operator, gauge natural bundle, affine connection.

This paper is in final form and no version of it will be submitted for publication elsewhere.

2. Gauge natural bundle of Cartan connections

Every principal connection can be interpreted as a section Γ of a fibred manifold J^{1}(P′) → P′ (with the projection on the target) invariant with respect to the right action of G on J^{1}(P′) (it is induced by the principal action of G on P′). Then the structure of fibred manifold can be introduced on the set QP′ = J^{1}(P′)/G and this construction gives rise to a functor Q : PB_{m}(G) → FM, the bundle of principal connections (see [KMS, 17.4.]).

Analogously we can study Cartan connections. There is a one-to-one correspondence between Cartan connections on P and principal connections on P′ with certain properties. We claim that these connections are even all sections of a subbundle of QP′. Thus Cartan connections can be viewed as sections of a subbundle of the bundle functor Q( ×_{B} G). We describe the correspondence in more detail.

Firstly we describe the correspondence between Cartan connections on P and principal connections on P′. A Cartan connection on P, that is a form ω ∈ Ω(P; 𝔤), can be spread as a pseudotensorial form of the type (Ad, 𝔤) (KN, II.5) over the whole P′ (P ⊂ P′, see above)

ω_{(u,g)} = Ad(g^{−1})π_{P}^{∗} ω + π_{G}^{∗}ω_{G},

where π_{p} : P × G → P and π_{G} : P × G → G are canonical projections, and ω_{G} is the left Maurer-Cartan form on G. Precisely speaking, we extended ω to a form on P ×G, but this extension is a pull-up of a connection form of a principal connection on P′ = P ×_{B} G. This construction gives rise to principal connections on P′, whose horizontal bundle doesn’t meet the tangent bundle of P viewed as the subbundle of P′ (see [S, Appendix A] for details). We will call this condition the condition †. Converselly the pull-back of a connection form of a principal connection which satisfies the condition †, under the canonical inclusion i : P → P ×_{B} G is a Cartan connection on P with values in 𝔤. These constructions are inverse to each other.

From now on we will use the term ”Cartan connection” also in the sense of a principal connection on P′ satisfying the condition †.

The condition † is obviously preserved by the action of PB_{m}(B)-morphisms: there is the embedding of PB_{m}(B) into PB_{m}(G) (see above) and thus we can consider any PB_{m}(B)-morphism as a PB_{m}(G)-morphism. The image of the horizontal bundle of the given connection under the PB_{m}(G)-morphism defines then the resulting connection. Since every PB_{m}(G)-morphism is a local diffeomorphism, its tangent mapping is an isomorphism in every point of P′, and consequently the new connection also satisfies the condition † (because of dimensions, the horizontal space of a Cartan connection in any point u ∈ P and the tangent space of P at u generate the tangent space of P′ at u, so must their images under any isomorphism and again because of dimensions their intersection has to be the zero vector only).

It makes sense to ask, whether a given principal connection satisfies the condition † in a given point of P′, that is if an element of QP satisfies the condition †. If so, we will call this element an element of Cartan connection. The action of PB_{m}(B)-morphisms on Cartan connections (i.e. principal connections with special properties) is given pointwise and we can as well consider the action just on the elements of Cartan connections. We have just proved, that these elements (i.e. values of Cartan connections in a given point of P′) are closed under the action of PB_{m}(B) morphisms, moreover they form an open subset of QP (because of the nature of the condition †). Thus these elements form a gauge natural subbundle of the gauge natural bundle Q( ×_{B} G) over the category PB_{m}(B). We will call this subbundle the bundle of Cartan connections C. That means the bundle of Cartan connections is a functor C : PB_{m}(B) → FM on the category of principal B-bundles over m-dimensional manifolds to the category of fibred manifolds. Then Cartan connections are sections of the bundle of Cartan connections.

Now we can study natural operators on Cartan connections as gauge natural invariants.

3. Natural operators on gauge natural bundles

In this paper we deal with the concept of a natural operator stated in [KMS]. Natural transformations (zero order natural operators) between two r-th or-

der gauge natural bundles correspond to W_{m}^{r}(B)-equivariant maps between their standard fibres (where W_{m}^{r}(B) is the group of r-jets of local isomorphisms of principal B-bundles). To get the higher order operators we have to work with the corresponding jet prolongation of the source bundle (see [KMS]). We are going to study first order operators on the first order gauge natural bundle, thus we will study the action of W_{m}^{2}(B) on the standard fiber of the first jet prolongation of the bundle of Cartan connections. W_{m}^{2}(B) is canonically embedded into W_{m}^{2}(G) and the desired action is the restriction of the action of W_{m}^{2}(G) on the standard fibre J_{0}^{1}(Q(ℝ^{m} × G)) of the first jet prolongation of the bundle of principal connections on ℝ^{m} × G (the standard fibre of J^{1}(Q(P′)) is isomorphic to the standard fibre of J^{1}(Q(ℝ^{m} × G)) for any principal bundle P′).

The following two paragraphs are quite technical ones.

4. The action of morphisms on principal connections

In this paragraph we will present the calculations made in [KMS, Section 52]: we will compute the action of PB_{m}(G)-morphisms on the standard fibre of the bundle of principal connections. Knowing this action we will be able then to compute in the following paragraph the action of PB_{m}(G)-morphisms on the standard fibre of Cartan connections quite easily.

For the standard fibre S of the bundle of principal connections we use identification S = J_{0}^{1}(ℝ^{m} ×G)/G ≃ J_{0}^{1}(ℝ^{m}, G)_{e} ≃ 𝔤⊗ℝ^{m∗}. In this identification we take a canonical representative Γ(x, e) ∈ ℝ^{m∗} ⊗ 𝔤 to represent a connection Γ in the fibre over x. In cordinates we get Γ_{i}^{p}, the Christoffel symbols (for a fixed base e_{p} in 𝔤); then the connection form ω of the connection Γ is related with the coordinates Γ_{i}^{p} in the standard fibre via the left Maurer-Cartan form ω_{G} of G by the formula Γ_{i}^{p}dx^{i} = ω_{G}(e) − ω(x, e).

Now we describe coordinates on W_{m}^{2}(G): for PB_{m}(G)-isomorphism Φ : ℝ^{m} × G → ℝ^{m} × G; Φ(x, g) = (f(x), φ(x) · g), f(0) = 0, φ : ℝ^{m} → G, φ(x) = Φ(x, e), the coresponding element in W_{m}^{2}(G) has the coordinates a = φ(0) ∈ G, (a_{i}^{p}, a_{ij}^{p}) ∼ j_{0}^{2}(a^{−1} · φ(x)) ∈ 𝔤 ⊗ ℝ^{m∗} ⊕ 𝔤 ⊗ S^{2}R^{m∗}, (a_{j}^{i}, a_{jk}^{i}) ∼ j_{0}^{2}f ∈ G_{m}^{2}. The coordinates on W_{m}^{1}(G) are then only the first order partial derivatives terms: (a, a_{j}^{i}, a_{i}^{p}).

The action of Φ on Γ(x) ≃ j_{x}^{1}s ∈ Q(ℝ^{m} ×G), is given by the formula QΦ(j_{x}^{1}s) = j_{f(x)}^{1}(y ↦ (ρ^{(φ(x))^{−1}} ∘ Φ ∘ s ∘ f^{−1}(y))) = j_{f(x)}^{1}(y ↦ (conj(φ(x)) ∘ μ ∘ (λ^{(φ(x))^{−1}} ∘ φ, s) ∘ f^{−1}(y))), where ρ^{a} is the right multiplication by an element a in G, λ is the left multiplication by a in G, and μ is the multiplication in G. This yields then in coordinates

\overline{Γ}_{i}^{p}(f(x)) = A_{q}^{p}(φ(x))(Γ_{j}^{q}(x) + a_{j}^{q}(x))a_{i}^{j},

where A_{q}^{p}(a) is a coordinate expression of the adjoint representation of G, a_{i}^{j} is the inverse matrix to a_{i}^{j}, that’s a_{j}^{i} ∼ j_{0}^{1}(f^{−1}). Especially for Γ(0) ∼ j_{0}^{1}(s) ∈ S we get

\overline{Γ}_{i}^{p} = A_{q}^{p}(a)(Γ_{j}^{q} + a_{j}^{q})a_{i}^{j}.

On J_{0}^{1}(Q(ℝ^{m} × G)) we introduce coordinates Γ_{j}^{p} and Γ_{jk}^{p} = ∂Γ_{j}^{p}/∂x^{k}.

The action of W_{m}^{2}(G) on J_{0}^{1}(Q(ℝ^{m} × G)) is then given by the formula (j_{0}^{2}Φ)(Γ) = j_{0}^{1}(QΦ ∘ Γ ∘ f^{−1}) ∼ j_{0}^{1}(QΦ ∘ j^{1}s ∘ f^{−1}). In coordinates we get

\overline{Γ}_{ij}^{p} = \frac{∂\overline{Γ}_{i}^{p}}{∂x^{j}} = \frac{∂}{∂x^{j}} (A_{q}^{p}(φ ∘ f^{−1}(x))(Γ_{k}^{q}(f^{−1}(x)) + a_{k}^{q}(f^{−1}(x)))a_{i}^{k})
= A_{q}^{p}(a)Γ_{kl}^{q}a_{i}^{k}a_{j}^{l} + A_{q}^{p}(a)a_{kl}^{q}a_{i}^{k}a_{j}^{l} + D_{qr}^{p}(a)Γ_{k}^{q}a_{l}^{r}a_{i}^{k}a_{j}^{l}
+ E_{qr}^{p}(a)a_{k}^{q}a_{l}^{r}a_{i}^{k}a_{j}^{l} + A_{q}^{p}(a)(Γ_{k}^{q} + a_{k}^{q})a_{ij}^{k},

where D_{qr}^{p} and E_{qr}^{p} are some function on G, which we won’t need, cf. [KMS].

5. Action of morphisms on Cartan connections

As promissed we will use the preceeding computation to express the action of PB_{m}(G) morphisms on the standard fibre of Cartan connections.

The action of W_{m}^{2}(B) on the J_{0}^{1}(C(ℝ^{m}×B)) is then given by the same formulae; what we are left with to do is to describe the restriction of J_{0}^{1}(Q(ℝ^{m} × G)) to J_{0}^{1}(C(ℝ^{m} × B)) and the embedding of W_{m}^{2}(B) into W_{m}^{2}(G) in coordinates.

5.1. Notation. We can choose a base e_{p} of 𝔤 in such a way, that e_{p}, p ≤ dim 𝔟 form a base of 𝔟, and we shall write (e_{p}) = (e_{s}, e_{t}), 1 ≤ s ≤ dim 𝔟, dim 𝔟 < t ≤ dim 𝔤, where (e_{s}) is a base of 𝔟. Then (e_{t}) represents a base in 𝔤/𝔟. Correspondingly for a vector v ∈ 𝔤 the ^{s} coordinates determine its 𝔟 part and ^{t} coordinates its 𝔤/𝔟 part.

The horizontal bundle HP′ of a principal connection on P′ corresponding to a Cartan connection satisfies the condition †, that is T P ∩ HP′ = P × {0}. This means that vectors Γ_{i}^{t}, i = 1, …,m are linearly independent and consequently the matrix Γ_{i}^{t} is regular. The coordinates on T := J_{0}^{1}(C(ℝ^{m} × B)) are then Γ_{i}^{p}, Γ_{ij}^{p}, with Γ_{i}^{t} a regular matrix and Γ_{ij}^{p} arbitrary.

The coordinates corresponding to the embedding of W_{m}^{2}(B) into W_{m}^{2}(G) are (a, a_{i}^{s}, a_{i}^{t} = 0, a_{ij}^{s}, a_{ij}^{t} = 0, a_{j}^{i}, a_{jk}^{i}). Notice, that the action of W_{m}^{1}(B) on the standard fibre of Q(P) really preserves the standard fibre of C(P) (it is exactly what it should do): it preserves the regularity of Γ_{i}^{t}.

6. The orbit reduction

Now we exploit the orbit reduction theorem (cf. [KMS, Section 28]) to simplify the action on the standard fibre. To be able to do this we introduce new coordinates on J_{0}^{1}C(ℝ^{m} × B). These new coordinates are Γ_{i}^{p}, Γ_{(ij)}^{p} (symetrizations of derivations of Christoffel symbols),

(2)  R_{ij}^{p} = Γ_{[ij]}^{p} + c_{qr}^{p}Γ_{i}^{q}Γ_{j}^{r},

where c_{qr}^{p} are the structure constants of the group G. Recall that 1 ≤ s ≤ dim 𝔟, dim 𝔟 < t, u ≤ dim 𝔤. R_{ij}^{p} coordinates are just the coordinate expression of the curvature tensor.

Now the kernel of the projection W_{m}^{2}B → G_{m}^{1} × B, (with coordinates a_{i}^{p}, a_{ij}^{p} – symmetric, a_{j}^{i} = δ_{j}^{i} a = e, a_{jk}^{i}) acts on these in the following way:

\overline{Γ}_{i}^{s} = Γ_{i}^{s} + a_{i}^{s}
\overline{Γ}_{i}^{t} = Γ_{i}^{t}
\overline{Γ}_{(ij)}^{s} = Γ_{(ij)}^{s} + a_{ij}^{s} + D_{qr}^{s}(e)Γ_{k}^{q}a_{l}^{r}a_{(i}^{k}a_{j)}^{l} + E_{qr}^{s}(e)a_{k}^{q}a_{l}^{r}a_{(i}^{k}a_{j)}^{l} + (Γ_{k}^{s} + a_{k}^{s})a_{ij}^{k}
\overline{Γ}_{(ij)}^{t} = Γ_{(ij)}^{t} + D_{qr}^{t}(e)Γ_{k}^{q}a_{l}^{r}a_{(i}^{k}a_{j)}^{l} + E_{qr}^{t}(e)a_{k}^{q}a_{l}^{r}a_{(i}^{k}a_{j)}^{l} + Γ_{k}^{t}a_{ij}^{k}
\overline{R}_{ij}^{p} = R_{ij}^{p}

The last equality results from the substance of the curvature tensor (the curvature is a natural operator). We can see that the kernel acts on Γ_{i}^{s}, Γ_{(ij)}^{p} transitively, and on Γ_{i}^{t} and R_{ij}^{p} trivially: it is obvious, that it acts transitively on Γ_{i}^{s}; as for Γ_{(ij)}^{t} and Γ_{(ij)}^{s} we can ”adjust” Γ_{ij}^{t} by the term Γ_{k}^{t}a_{ij}^{k} (Γ_{k}^{t} is a regular matrix and a_{ij}^{k} arbitrary) and then we can choose freely a_{ij}^{s} and thus we get a transitive action on Γ_{(ij)}^{s}. That means the action of W_{m}^{2}B on T = J_{0}^{1}C(ℝ^{m} × B) can be factorized through the action of G_{m}^{1} × B just on coordinates Γ_{i}^{t}, R_{ij}^{p}.

We conclude that every W_{m}^{2}(B)-map from T to (G_{m}^{1} ×B)-space Z (every (G_{m}^{1} × B)-space can be viewed as a W_{m}^{2}(B)-space – we consider the action of the kernel of the projection W_{m}^{2}B → G_{m}^{1} × B to be trivial) can be factorized through G_{m}^{1} × Bmap defined just on coordinates Γ_{i}^{t} and R_{ij}^{p} (with the above action) to Z.

If we denote by α : T → 𝔤/𝔟 ⊗ ℝ^{m∗} the projection (Γ_{i}^{p}, Γ_{ij}^{p}) ↦ (Γ_{i}^{t}) and by β : T → 𝔟 ⊗ Λ^{2}ℝ^{m∗} the map (2), we can summarize:

6.1. Proposition. Let Z be a (G_{m}^{1} × B)-space. Then for every W_{m}^{2}B-map f : T → Z there exists a unique (G_{m}^{1} × B)-map g : 𝔤/𝔟 ⊗ ℝ^{m∗} ⊕ 𝔤 ⊗ Λ^{2}ℝ^{m∗} → Z satisfying f = g ∘ (α, β).

7. The main theorem and its application

We can interpret Proposition 6.1. in terms of natural operators: The first order gauge natural operators on the bundle of Cartan connections with target at the gauge natural bundles, where just the subgroup G_{m}^{1} × B of W_{m}^{2}B acts nontrivially (the value of the bundle (i.e. of the functor) on PB_{m}(B)-morphism in a given point depends only on 1-jet of base mapping of the given morphism and only on the value of the given mapping in the fibre – the bundles of the order (1,0), see [KMS, Section 52]), depends only at the curvature form of the given connection and 𝔤/𝔟-part of Christoffel symbols. For a given Cartan connection on P, the curvature form is so called tensorial 2-form of the type (𝔤, Ad), i.e. it is a horizontal 2-form and r^{a∗}Ω = Ad^{−1}(a)Ω, thus it can be regarded as a section of the bundle LP′⊗Λ^{2}T (BP)^{∗}, where P′ is the principal G-bundle defined in Section 2, B is the base functor, and LP is the asociated bundle LP := P′ ×_{Ad} 𝔤.

We get the following result:

7.1. Theorem. Let K be a gauge natural bundle of order (1, 0). Then for every first order gauge natural operator A : C → K there exists a unique natural transformation O : (L ⊗ Λ^{2}T^{∗}B) ⊕ (𝔤/𝔟 ⊗ T^{∗}B) → K satisfying A = O ∘ (E, F), where E : C → L ⊗ Λ^{2}T^{∗}B is the curvature operator, and F : C → 𝔤/𝔟 ⊗ T^{∗}B is the 𝔤/𝔟-part of a Cartan connection (a 𝔤-valued form).

7.2. Relations to natural operators on affine connections.

Now we are going to apply Theorem 7.1 in the situation, where B ⊂ GL(n, ℝ) and G = B ⋊ℝ^{m} to get the classical result on standard operators on affine connections. Let ρ be a Cartan connection with values in 𝔤 on the principal B-bundle P and π : P → BP the projection on a base manifold. Such a connection can be split to ρ = ω + φ according to the splitting 𝔤 = 𝔟 + ℝ^{m}. The 𝔟 part is just a principal connection on P, the ℝ^{m} part represents a non-degenerated tensorial 1-form of the type (GL(n, ℝ), ℝ^{m}) on P (see [KN]), that is it is a horizontal form and it is right equivariant with respect to the actions of B – see the second condition in the definition of the Cartan connection). We will write φ_{u} : T_{π(u)}M → ℝ^{m}. This form identifies P with a subbundle of the bundle of linear frames P^{1}M, namely we have an injection i : P → P^{1}M, i(u) = φ_{u}^{−1} (we consider a linear frame at x ∈ M to be a linear mapping from ℝ^{m} to T_{x}M). Then i is a principal bundle morphism:

i(u · A) = φ_{u·A}^{−1} = (A^{−1} · φ_{u})^{−1} = φ_{u}^{−1} ∘ A = i(u) · A,

for A ∈ B. Then the push out of the form φ under the morphism i is just the canonical soldering form θ on i(P) ⊂ P^{1}M:

i_{∗}φ(X_{E}) = φ_{i^{−1}(E)}π(X_{E}) = E^{−1}(π(X_{E})) = θ(X_{E}),

for a linear frame E ∈ i(P).

Let us explain, what we mean by an affine connection in this paper. An affine connection will be a ”classical” affine connection a (a is 1-form on P^{1}M, a = ω+θ, where ω is a linear connection on P^{1}M and θ is the soldering form, [KN, Chapter VI]), whose restriction to P ⊂ P^{1}M has values in 𝔤 = 𝔟 + ℝ^{m}. In other words it is a principal connection belonging to a B-structure – the realization of the Bstructure as a subbundle of P^{1}M is enabled by the soldering form. Equivalently we can say, that an affine connection is a Cartan connection a on a B-subbundle P of the bundle of linear frames and a = ω + θ, where θ is the restriction of the soldering form θ on P^{1}M to the subbundle P.

We will show, that any first order gauge natural operator on Cartan connections on P of the discussed type (i.e. operator invariant under the action of PB_{m}(B) morphisms) can be restricted to a first order natural operator on affine connections (i.e. operator invariant under the action of morphisms of the form P^{1}f, f : BP → BP) and converselly any first order natural operator on affine connections can be extended to a first order gauge natural operator on Cartan connections.

It is quite obvious, that each gauge natural operator on Cartan connections with values in 𝔤 defines a natural operator on affine connections: restricting the given operator just to affine connections (Cartan connections on a subbundle of P^{1}M with some extra properties) we obtain a natural operator on affine connections. Such an operator is then equivariant with respect to the action of P^{1}(f) morphisms f : ℝ^{m} → ℝ^{m} (the action of morphism on sections of source or target bundles is given by pull-backs or equivalently by push-outs), because the original operator was equivariant to any principle bundle morphism.

Conversely, given a first order natural operator on affine connections, we can extend it to an operator over the whole PB_{m}(B). Since any first order gauge natural operator on Cartan connections is determined by its values on the standard fibre J_{0}^{1}C(ℝ^{m} ×B), we will carry on our considerations on the bundle ℝ^{m} ×B. Let P = ℝ^{m} × B be a principal bundle with structure group B and we have chosen a tensorial form θ_{0} of the type (GL(n, ℝ), ℝ^{m}) which identifies P with a subbundle of P^{1}(BP). Further let Z be a gauge natural bundle of order (1,0) and D : A → Z be a natural operator on affine connections. Then we define a natural operator \overline{D} : C → Z. Let c = ω + θ_{1} be a splitting of a Cartan connection on P. Now the forms θ_{0} and θ_{1} are related in the following way: their values at (0, b), b ∈ B are given by values at (0, e). If E ∈ GL(n, ℝ) is a matrix representing θ_{0} at (0, e) and F ∈ GL(n, ℝ) a matrix representing θ_{1} in (0, e) then F = EE^{−1}F. If we choose a PB_{m}(B) morphism φ = (φ_{0}, const_{e}) with j_{(0,e)}^{1}φ_{0} = E^{−1}F ∈ G_{m}^{1}, then (1) yields θ_{1}(0) = φ^{∗}θ_{0}(0) and we define (for a value of an operator \overline{D} on a 1-jet of connection c at 0 ∈ ℝ^{m}; we denote the corresponding element in J_{0}^{1}C(ℝ^{m} × B) by the same letter c)

\overline{D}(c) = \overline{D}(ω+θ_{1}) = \overline{D}((φ^{∗}(φ_{∗}ω)+φ^{∗}θ_{0}) = φ^{∗}\overline{D}(φ_{∗}ω+θ_{0}) = φ^{∗}i^{∗}D(i_{∗}(φ_{∗}ω+θ_{0})).

Since our choice of 1-jet at (0, e) of the morphism φ is uniquely determined by θ_{1} and θ_{0}, this definition is correct. Now we can decompose any morphism ψ : ℝ^{m} × B → ℝ^{m} × B to (ψ_{0}, \overline{ψ}) (see paragraph 4) and we can write for a 1-jet of the Cartan connection c in 0 ∈ ℝ^{m}, c = ω + θ_{1}:

\overline{D}(ψ^{∗}c) =\overline{D}(\overline{φ} ∘ ψ_{0} ∘ (\overline{ψ} ∘ \overline{φ}^{−1}))^{∗}c)
= \overline{D}((\overline{ψ} ∘ \overline{φ}^{−1})^{∗} ∘ ψ_{0}^{∗} ∘ \overline{φ}^{∗}(c))
= (\overline{ψ} ∘ \overline{φ}^{−1})^{∗}D(ψ_{0}^{∗} ∘ \overline{φ}^{∗}(c))
= (\overline{ψ} ∘ \overline{φ}^{−1})^{∗} ∘ ψ_{0}^{∗}D(\overline{φ}^{∗}(c))
= (\overline{ψ} ∘ \overline{φ}^{−1})^{∗} ∘ ψ_{0}^{∗} ∘ \overline{φ}^{∗}\overline{D}(c))
= ψ^{∗}\overline{D}(c),

which proves \overline{D} to be a gauge natural operator.

Taking a form θ′_{0} to identify ℝ^{m} × B with P^{1}(ℝ^{m}) (we will write i′ for the corresponding isomorphism) we would get the same operator: let θ′_{0} = η^{∗}θ_{0}, θ_{1} = ψ^{∗}θ′_{0}, θ_{1} = φ^{∗}θ_{0}, that is η ∘ ψ = φ. Then

\overline{D}(c) = \overline{D}(ω + θ_{1}) = \overline{D}((ψ^{∗}(φ_{∗}ω) + ψ^{∗}θ′_{0})
= ψ^{∗}\overline{D}(ψ_{∗}ω + θ′_{0}) = ψ^{∗}(i′)^{∗}D(i′_{∗}(ψ_{∗}ω + θ′_{0}))
= (η^{−1} ∘ φ)^{∗}(i ∘ η)^{∗}D((i ∘ η)_{∗}(ψ_{∗}ω + θ′_{0}))
= φ^{∗} ∘ (η^{−1})^{∗}η^{∗}i^{∗}D(i_{∗} ∘ (η_{∗}ψ_{∗}ω + η_{∗}θ′_{0}))
= φ^{∗}i^{∗}D(i_{∗}(φ_{∗}ω + θ_{0})).

That means we have proved:

7.3. Theorem. There is a one-to-one correspondence between natural operators on affine connections and gauge natural operators on Cartan connections.

7.4. Example. We can apply the Theorems 7.1 and 7.3 to deduce, that natural operators on affine connections (that is Cartan connections on bundle of linear frames with values in 𝔞𝔣(n, ℝ)) can be factorized through curvature form of an affine connection, that is of the curvature and torsion form of a corresponding linear connection. It is namely obvious, that natural operators on affine connections can’t depend on the soldering form θ, since they should be invariant to the action of morphisms preserving θ, and θ plays a role of a ”passive invariant”: it is invariant by definition but it doesn’t come up explicitly in any ”real” invariant.

References

[KN] Kobayashi, S., Nomizu, K., Foundations of differential geometry, Interscience Publishers, New York, London, 1963.

[KMS] Kolář, I., Michor, P., Slovák, J., Natural operations in differential geometry, Springer, Berlin New York, 1993.

[S] Sharpe, R. W., Differential geometry, Springer, New York Berlin Heidelberg, 1997.

Mathematics Institute, Academy of Science of the Czech Republic Žižkova 22, 616 62 Brno, CZECH REPUBLIC E-mail: naca@math.muni.cz